Suppose that the function y = f(x) is increasing on the interval (4, 7). Over what interval is the graph of
y = f(x + 8) increasing?

Answer by Theo(13342)

y = f(x+8) is shifted to the left 8 units.

therefore, if f(x) is increasing in the interval of (4,7), then y = f(x + 8) should be increasing in the interval of (-4,-1).

you should be able to take any function and show that this is true.

for example:

let y = f(x) = x^2 - 10

when x = 4, y = 6
when x = 7, y = 39

let y = f(x+8) = (x+8)^2 - 10

when x = -4, y = 6
when x = -1, y = 39

note that -4 + 8 = 4 and -1 + 8 = 7.

so f(4) = 6 and f(-4+8) = f(4).